OpenStudy (vortish):

completely factor the polynomial 3x^3+9x^2=3X +9

OpenStudy (campbell_st):

well on the left hand side of the equation you can factor out \[3x^2\] so you get \[3x^2(x + 3)\] on the right hand side of the equation you can factor out 3 \[3(x + 3)\] so you have \[3x^2(x + 3) = 3(x + 3)\] there is still a bit to do but hopefully this will help you get to the answer

OpenStudy (vortish):

IT does thanks

OpenStudy (vortish):

now im stuck

OpenStudy (campbell_st):

ok.. so rewrite the equation as 3x^2 (x + 3) -3(x +3) = 0 so you get (x +3)(3x^2 - 3) = 0 factor out the 3 in the 2nd set of brackets and you have 3(x + 3)(x^2 -1) = 0 now the 2nd set of brackets in the difference of 2 squares which you'll need to factor.. hope it makes sense.

OpenStudy (anonymous):

so now u factor that last bracket :) make it into ( ) ( ) form :)